A solution contains 4 grams of sodium hydroxide in 250ml of solution . If Na=23, O=16, H=1 and
1cm3
of water weighs 1g, what isthe concentration of the solution in grams per
dm3
?
12
14
16
18
1
Expert's answer
2015-10-22T08:59:55-0400
A solution contains 4 grams of sodium hydroxide in 250ml of solution.
1 cm3 of H2O = 0.01 dm3 H2O
m(NaOH) = 4 g
V(Solution) = 250 mL = 0.25 dm3
w(NaOH) = m(NaOH)/ V(Solution) = 4/0.25 = 16 g/dm3

Answer : The concentration of the solution is 16 g/dm3 .